OpenStudy (anonymous):

solve the system by the elimination method 5x+5y=-7 7x-3y=13 as an ordered pair of fractions

OpenStudy (anonymous):

\[(\frac{22}{25},-\frac{57}{25})\]

OpenStudy (anonymous):

wanna cheat? i will show you snap way

OpenStudy (anonymous):

thats not it

OpenStudy (anonymous):

to teach me a lesson?

OpenStudy (anonymous):

well actually it is

OpenStudy (anonymous):

as a fraction ordered pair?

OpenStudy (anonymous):

the answer i sent was right. really. i am not trying to trick you

OpenStudy (anonymous):

ordered pair means \[(x,y)\]

OpenStudy (anonymous):

both x and y can be fractions. they do not have to be integers

OpenStudy (anonymous):

ok, i will type in your answer

OpenStudy (anonymous):

on line class? i will be happy to show you how to do this the snap way if you like

OpenStudy (anonymous):

actually two snap ways

OpenStudy (anonymous):

it was incorrect

OpenStudy (anonymous):

well sorry but it is correct. i am assuming that you want the intersection of the two lines given by the equations \[5x+5y=-7\] \[7x-3y=13\] yes?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

aren't those the equations written above?

OpenStudy (anonymous):

solve the system by the elimination method

OpenStudy (anonymous):

which is a method for finding the intersection of the two lines. i.e. fining x and y that satisfy both equations.

OpenStudy (anonymous):

Wait, I will verify

OpenStudy (anonymous):

yes bliss please do

OpenStudy (anonymous):

i was told my answer was wrong, and i am not going to fret too much but i am fairly certain they are right

OpenStudy (anonymous):

I get 44/50 for x, that's the un-simplified version, so y must be -114/50

OpenStudy (anonymous):

so the solution is (44/50, -114/50)

OpenStudy (anonymous):

i think 44/50=22/25 yes?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That's why I said un-simplified

OpenStudy (anonymous):

Maybe her program likes un-simplified answers

OpenStudy (anonymous):

the correct answer was (22/25,-57/25)

OpenStudy (anonymous):

You should post the complete question, like maybe if it wants decimal form, simplified or un-simplified

OpenStudy (anonymous):

Satellite had it right then

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

I was just giving you an alternate solution

OpenStudy (anonymous):

@cali that is exactly what i wrote!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

look above that it the exact answer i sent you